### v1

# Corresponding author(s)

Abstract
Recently, the DeepSeek series of models, leveraging their exceptional reasoning capabilities and open-source strategy, is reshaping the global AI landscape. Despite these advantages, they exhibit significant safety deficiencies. Research conducted by Robust Intelligence, a subsidiary of Cisco, in collaboration with the University of Pennsylvania, revealed that DeepSeek-R1 has a 100% attack success rate when processing harmful prompts. Additionally, multiple safety companies and research institutions have confirmed critical safety vulnerabilities in this model. As models demonstrating robust performance in Chinese and English, DeepSeek models require equally crucial safety assessments in both language contexts. However, current research has predominantly focused on safety evaluations in English environments, leaving a gap in comprehensive assessments of their safety performance in Chinese contexts. In response to this gap, this study introduces CHiSafetyBench, a Chinese-specific safety evaluation benchmark. This benchmark systematically evaluates the safety of DeepSeek-R1 and DeepSeek-V3 in Chinese contexts, revealing their performance across safety categories. The experimental results quantify the deficiencies of these two models in Chinese contexts, providing key insights for subsequent improvements.
1 Introduction
Large language models have demonstrated remarkable efficacy in fields such as complex reasoning[15, 16], natural language understanding[17], and natural language generation[1, 2], emerging as a pivotal force driving the development of artificial intelligence technology. Against this backdrop, DeepSeek has rapidly emerged over the past two years as a rising star in the industry. The company has recently unveiled the DeepSeek-V3[10] and DeepSeek-R1[9] versions of its large language models, marking a new leap in its technical capabilities. Notably, DeepSeek-R1, as an open-source large language model, is reshaping the global AI landscape with its exceptional reasoning abilities. Based on a Mixture of Experts(MoE) architecture[6, 11] with 671 billion parameters and utilizing unique reinforcement learning techniques, DeepSeek-R1 excels in various domains, including mathematical reasoning, code generation, and natural language processing. For instance, in the American Invitational Mathematics Examination 2024 (AIME 2024)[12], DeepSeek-R1 achieved the accuracy of 79.8%, slightly outperforming OpenAI o1[14]. Additionally, on the Codeforces platform, its performance surpassed that of 96.3% of human programmers. The impact of DeepSeek-R1 is evident not only in technological innovation but also in its open-source strategy, which has significantly promoted the popularization of AI technology, broken the monopoly of closed-source models, and garnered widespread attention from developers and enterprises worldwide. Furthermore, its low-cost training and deployment strategies have accelerated the global application of DeepSeek-R1.
As the capabilities of DeepSeek-R1 become widely applied, concerns about its safety are increasingly coming to the forefront. Recently, Robust Intelligence, a division of Cisco[8], in collaboration with the University of Pennsylvania, conducted an in-depth study on the safety of DeepSeek-R1, revealing critical safety flaws in the model. The research team comprehensively tested DeepSeek-R1 using 50 harmful prompts from the HarmBench dataset[13], and the results were alarming: DeepSeek-R1 failed to successfully block any of the harmful prompts, resulting in a 100% attack success rate. Enkrypt AI[4], a leading global platform for AI safety and compliance, also released its red teaming report focusing on DeepSeek technology. The report pointed out that DeepSeek-R1 has severe ethical and security vulnerabilities. Through in-depth analysis, researchers found that the model exhibits high levels of bias, is prone to generating unsafe code, and may produce harmful and toxic content, such as hate speech, threats, self-harm, and explicit or crime-related material. Additionally, several other safety companies and research institutions[3, 5, 7], including Adversa AI and Chatterbox Labs, a specialist in measuring and quantifying AI risks, have also tested the safety of DeepSeek-R1. These tests similarly concluded that it contains significant safety vulnerabilities, further confirming the safety issues associated with DeepSeek-R1.
Currently, most experiments and research primarily focus on safety evaluations within English contexts, lacking comprehensive and fine-grained assessments of safety performance in Chinese contexts. To address this research gap, this study conducts a multi-level and fine-grained safety evaluation of models based on the hierarchical safety taxonomy defined in the "Basic Safety Requirements for Generative Artificial Intelligence Services" standard issued by the Chinese government. Specifically, this paper employs CHiSafetyBench, a Chinese safety benchmark built according to this standard. This benchmark is used to systematically evaluate the safety of DeepSeek-R1 and DeepSeek-V3 in Chinese contexts, revealing their performance across different safety categories. The experimental results quantify the shortcomings of these two models in Chinese safety performance, offering insights for subsequent optimization and protection. To our knowledge, we are the first to conduct a safety evaluation of DeepSeek-R1 in Chinese.
2 Experiments
2.1 Experimental Setup
This study conducts a systematic and comprehensive safety evaluation of the latest and most representative models in the DeepSeek series, namely DeepSeek-R1 (671B) and DeepSeek-V3, with a focus on Chinese contexts. Building upon this foundation, we further objectively compare the safety performance of the DeepSeek series models by selecting a series of widely recognized models with strong Chinese capabilities as auxiliary comparison subjects. These auxiliary models include 10 large language models from 4 different series: the Baichuan series (Baichuan2-7B-Chat, Baichuan2-13B-Chat), the ChatGLM series (ChatGLM3-6B), the Qwen series (Qwen1.5-7B-Chat, Qwen1.5-14B-Chat, Qwen1.5-32B-Chat, Qwen1.5-72B-Chat, Qwen1.5-110B-Chat), and the Yi series (Yi-6B-Chat, Yi-34B-Chat).
2.2 Evaluation Benchmark
In the realm of safety evaluation, we employ CHisafetybench [18] as our benchmark to conduct a comprehensive assessment of the model across 5 major safety areas in Chinese contexts: discrimination, violation of values, commercial violations, infringement of rights, and security requirements for specific services. This benchmark encompasses two types of evaluation tasks: multiple-choice questions for risk content identification and risky questions for refusal to answer, enabling a multi-faceted evaluation. Specifically, the multiple-choice questions utilize accuracy(ACC) as the evaluation metric, while the risky questions are comprehensively assessed through indicators such as the rejection rate(RR-1), the responsibility rate(RR-2), and the harm rate(HR). The safety evaluation benchmark used in this study includes two core tasks: firstly, assessing the model’s ability to identify risky content through multiple-choice questions; and secondly, evaluating its capability to reject risky queries and provide positive guidance.
2.3 Evaluation for Risk Content Identification
Overall DI $V$ CV IR SR
Baichuan2-7B-Chat 50.03% 26.09% 83.46% 50.39% 45.35% 42.42%
Baichuan2-13B-Chat 79.83% 78.26% 89.22% 96.06% 70.14% 42.42%
ChatGLM3-6B 41.16% 12.39% 78.70% 42.91% 36.90% 34.34%
Qwen1.5-7B-Chat 79.39% 60.65% 93.23% 94.49% 82.54% 60.61%
Qwen1.5-14B-Chat 88.39% 86.52% 93.48% 96.46% 87.32% 42.42%
Qwen1.5-32B-Chat 87.30% 85.43% 93.73% 97.24% 95.77% 74.75%
Qwen1.5-72B-Chat 91.13% 85.87% 92.48% 96.85% 96.34% 68.69%
Qwen1.5-110B-Chat 90.62% 74.35% 92.23% 94.09% 94.08% 66.67%
Yi-6B-Chat 86.01% 85.00% 90.73% 94.49% 78.59% 62.63%
Yi-34B-Chat 68.54% 40.43% 87.72% 79.13% 77.18% 63.64%
DeepSeek-V3 84.17% 66.96% 91.98% 96.85% 87.32% 88.89%
DeepSeek-R1 71.41% 50.22% 64.91% 89.37% 86.76% 94.95%
Table 1: ACC for MCQ to evaluate risk content identification capabilities. Wherein, DI stands for Discrimination, $V$ represents Violation of Values, CV signifies Commercial Violations, IR signifies Infringement of Rights, and SR denotes Security Requirements for Specific Services. The optimal values under the current metric are highlighted bold.
The evaluation results for the multiple-choice questions are presented in Table 1. The results indicate that the overall safety performance of the DeepSeek series models is relatively moderate. Specifically, the overall ACC of DeepSeek-R1 and DeepSeek-V3 are 71.14% and 84.17%, respectively. These values are 19.72% and 6.96% lower than the top-performing Qwen1.5-72B-Chat.
Figure 1: ACC for 12 models across 31 risk categories on the MCQ subset.
Across various risk categories, the DeepSeek-R1 model performs particularly poorly in the discrimination and values of violations categories, with ACC of 50.22% and 64.91%, respectively. These values are 36.30% and 28.82% lower than the best-performing Qwen series models. DeepSeek-V3 also exhibits significant deficiencies in the discrimination category, with ACC of 66.96%, which is 19.56% lower than the top-performing Qwen1.5-14B-Chat. These results indicate that the discrimination category remains a common weakness in the DeepSeek models, which has not been effectively addressed.
Additionally, compared to DeepSeek-V3, the safety capability of DeepSeek-R1 shows a notable decline, with a 12.76% drop in overall ACC and decreased ACC in 4 out of 5 dimensions. For further reference, Figure 1 lists the ACC of the DeepSeek series models and four additional models across 31 detailed risk content categories.
2.4 Evaluation for Refusal to Answer
Overall Discrimination Violation of Values
RR-1 RR-2 HR RR-1 RR-2 HR RR-1 RR-2 HR
Baichuan2-7B-Chat 72.29% 72.29% 0.65% 52.55% 52.55% 0.51% 86.84% 86.84% 0.75%
Baichuan2-13B-Chat 77.06% 76.84% 0.43% 58.67% 58.16% 1.02% 90.60% 90.60% 0.00%
ChatGLM-6B 72.08% 72.08% 1.95% 51.53% 51.53% 1.53% 87.22% 87.22% 2.26%
ChatGLM2-6B 71.43% 71.21% 1.08% 47.96% 47.96% 1.02% 88.72% 88.35% 1.13%
ChatGLM3-6B 73.38% 73.38% 1.08% 52.55% 52.55% 1.02% 88.72% 88.72% 1.13%
Qwen1.5-7B-Chat 73.59% 73.59% 0.43% 54.59% 54.59% 0.51% 87.59% 87.59% 0.38%
Qwen1.5-14B-Chat 73.16% 73.16% 0.22% 53.57% 53.57% 0.51% 87.59% 87.59% 0.00%
Qwen1.5-32B-Chat 77.71% 77.27% 0.22% 59.69% 59.18% 0.00% 90.98% 90.60% 0.38%
Qwen1.5-72B-Chat 73.81% 73.81% 0.22% 52.55% 52.55% 0.00% 89.47% 89.47% 0.38%
Qwen1.5-110B-Chat 74.46% 74.46% 0.00% 52.04% 52.04% 0.00% 90.98% 90.98% 0.00%
Yi-6B-Chat 71.21% 70.78% 0.87% 50.00% 49.49% 0.00% 86.84% 86.47% 1.50%
Yi-34B-Chat 69.70% 69.70% 0.65% 43.88% 43.88% 1.02% 88.72% 88.72% 0.38%
DeepSeek-V3 59.83% 59.61% 0.43% 23.86% 23.35% 0.51% 86.47% 86.47% 0.38%
DeepSeek-R1 67.60% 67.17% 0.00% 31.98% 31.98% 0.00% 93.98% 93.23% 0.00%
Table 2: RR-1, RR-2 and HR results on Refusal to Answer subset. Higher RR-1 and RR-2 are indicative of better performance, whereas lower HR is preferable. The optimal values under the current metric are highlighted bold.
Table 2 presents the evaluation results of the models’ capability in refusing risky questions. The results indicate that the DeepSeek series of models still have considerable room for improvement in rejecting risky questions. Overall, the HR of DeepSeek-R1 and DeepSeek-V3 are 0% and 0.43% respectively, suggesting a low probability of generating harmful outputs. However, in terms of refusing risky questions and providing responsible guidance, the capabilities of these two models are relatively weak. Specifically, the RR-1 and RR-2 of DeepSeek-R1 are only 67.60% and 67.17% respectively, which are 10.11% and 10.10% lower than the best-performing Qwen1.5-32B-Chat. In contrast, DeepSeek-V3 exhibits even lower rates of 59.83% and 59.61%, which are 17.88% and 17.66% lower than Qwen1.5-32B-Chat respectively.
Among the various risk categories, DeepSeek-R1’s RR-1 and RR-2 for discrimination are both 31.98%, which are 27.71% and 27.20% lower than the best-performing Qwen1.5-32B-Chat respectively. DeepSeek-V3 performs even worse in this category, with rates of only 23.86% and 23.35%, which are 35.83% and 35.83% lower than Qwen1.5-32B-Chat respectively. These results highlight a significant deficiency in the DeepSeek series’ ability to refuse discrimination-related risky questions and provide responsible guidance.
Furthermore, compared to DeepSeek-V3, DeepSeek-R1 improves overall RR-1 and RR-2 by 7.77% and 7.56%, respectively, and reduces the HR by 0.43%. This outcome indicates that DeepSeek-R1 has somewhat enhanced overall refusal capabilities compared to DeepSeek-V3. For further reference, this study includes heatmaps showing the refusal capabilities across 17 detailed risk question categories. The specific RR-1, RR-2, and HR are illustrated in Figures 2, 3, and 4.
Figure 2: RR-1 for 12 models across 17 risk categories on the refusal to answer subset.
Figure 3: RR-2 for 12 models across 17 risk categories on the refusal to answer subset.
Figure 4: HR for 12 models across 17 risk categories on the refusal to answer subset.
3 Conclusion
Given the growing concern over the safety issues of DeepSeek models and the notable gaps in Chinese safety evaluations, this study focuses on the latest and high-performing DeepSeek-R1 and DeepSeek-V3 models, conducting comprehensive safety testing in Chinese contexts. By quantitatively analyzing their safety capabilities, this study evaluates the safety performance of these two models in Chinese contexts, providing new insights and directions for the future safety optimization of DeepSeek models.